OpenStudy (anonymous):

Find integers that are upper and lower bounds for the real zeros of the polynomial. (Be sure the lower bound is the largest possible lower bound and the upper bound is the smallest possible upper bound.) P(x) = x3 − 3x2 + 4

OpenStudy (anonymous):

wait....it should be x^3-3x^2+4

OpenStudy (anonymous):

So they want the infimum and the supremum.

OpenStudy (anonymous):

if that is the upper and lower bounds, then yes

OpenStudy (anonymous):

Least upper bound and greatest lower bound.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

This is a calculus course, no?

OpenStudy (anonymous):

its college algebra

OpenStudy (anonymous):

Ah.

OpenStudy (anonymous):

That polynomial is factorable. If it had a rational zero it must be an integer that divides 4, so the options are\[\pm 1, \pm 2, \pm 4.\]You test these numbers and find that \[-1\]is a root of the polynomial and then use long division to get\[\frac{x^3 - 3x^2 + 4}{x -(-1)} = x^2 - 4x + 4 = (x - 2)^2\]so \[x^3 - 3x^2 + 4 = (x+1)(x-2)^2.\]Therefore, the roots are -1 and 2 (the second with multiplicity 2) and the lower and upper bounds for the real roots are -1 and 2 respectively.